OpenStudy (anonymous):

In the equation x – 4y = 2, for what value of x will y = 0?

OpenStudy (anonymous):

A. 2 B. 6 C. –2 D. 0

OpenStudy (anonymous):

x – 4y = 2, put y=0 we find x- 4*0=2 i.e. x-0=2 i.e. x=2 hence at y=0 the value of x is 2

OpenStudy (anonymous):

thanks